OpenStudy (anonymous):

find the area of circle with radius=17yd.use pi»3.14 the area is approximately

OpenStudy (anonymous):

use the formula of area of the circle

OpenStudy (anonymous):

its pi*r*r=3.14*17

OpenStudy (anonymous):

*17

OpenStudy (anonymous):

no its 17^2 which is 289

OpenStudy (anonymous):

and then times by 3.14 which is

OpenStudy (anonymous):

i think its 9076 yd2

OpenStudy (anonymous):

no 879

OpenStudy (anonymous):

no its 907.46 yd^2

OpenStudy (anonymous):

you right i mass up

OpenStudy (anonymous):

ya 907.46

OpenStudy (anonymous):

i got a new one

OpenStudy (anonymous):

16yd.use pi»3.14

OpenStudy (anonymous):

ok formula of area of circle is A=3.14 x r^2 A=3.14 x 16^2 16^2=256 A=3.14x256 A=803.84

OpenStudy (anonymous):

its 803.84

OpenStudy (anonymous):

yes 803.84 is the answer....